OpenStudy (anonymous):

find all the zeros to y = x^3+3x^2+2x

OpenStudy (gorv):

=x(x^2+3x+2) =x(x^2+2x+x+2) =x(x(x+2)+x+2) =x(x+2)(x+1) so zeros are x=0 x=-1 x=-2